OpenStudy (anonymous):

find the range 42 to (-93)

OpenStudy (anonymous):

42-(-93)=135

jimthompson5910 (jim_thompson5910):

Subtract the smaller number from the larger number 42 - (-93) = 42+93 = 135

OpenStudy (saifoo.khan):

C'mon Who'z gonna win??

OpenStudy (saifoo.khan):

Meera won!

jimthompson5910 (jim_thompson5910):

so the range is 135

OpenStudy (saifoo.khan):

but still Jim Rocks as well. :D

jimthompson5910 (jim_thompson5910):

lol thx

OpenStudy (anonymous):

I wanna steal you saifoo!!!

OpenStudy (saifoo.khan):

haha, why is that?

jimthompson5910 (jim_thompson5910):

not sure that's legal in all states....

OpenStudy (saifoo.khan):

http://youtu.be/vSEYrnUSi6I

OpenStudy (anonymous):

hahah he is adorable :P

OpenStudy (anonymous):

ikr! lol i got dibs yesterday =]

OpenStudy (anonymous):

haha what the heck -.-

OpenStudy (anonymous):

the youtube thing

OpenStudy (saifoo.khan):

haha, thanks for the complements. i just posted that idk why. LOL. @meeru, what does dibs stand for? :O

OpenStudy (anonymous):

its like i get it first...like when you see something you like you say dibs and that means you called it first and its yours =]

OpenStudy (saifoo.khan):

Oh, now i get it.. lol @ your comment. ;]